OpenStudy (anonymous):

find the area of sector of a circle whose radius is 6 cm and length of corresponding arc is 12 cm.

ganeshie8 (ganeshie8):

u may setup a proportion

ganeshie8 (ganeshie8):

\(\large \frac{12}{2\pi\times 6} = \frac{x}{\pi\times 6^2} \)

ganeshie8 (ganeshie8):

simplify a bit, cross multiply and solve \(x\)

OpenStudy (anonymous):

x is what?..here?

ganeshie8 (ganeshie8):

x is the required area of sector

OpenStudy (anonymous):

can we solve this by this formula \[\theta \div 360 * 2\pi r\]

OpenStudy (anonymous):

in this i need to find the \[\theta\]...how can i find that..and what is the use of corresponding arc?

ganeshie8 (ganeshie8):

you need to use : \(\theta \div 360 * \pi r^2 \)

ganeshie8 (ganeshie8):

Area of sector = \(\large \frac{\theta}{ 360 } \pi r^2\)

OpenStudy (anonymous):

yup

ganeshie8 (ganeshie8):

or in radians it wud be :- Area of sector \(\large \frac{1}{ 2} r^2\theta\)

ganeshie8 (ganeshie8):

you can find \(\theta \) in radians by using below :- \(\theta = arclength / radius\)

ganeshie8 (ganeshie8):

since we're given, arc length = 12 radius = 6, u can find \(\theta\)

ganeshie8 (ganeshie8):

\(\theta = 12/6 = 2\) radians

ganeshie8 (ganeshie8):

so, the area wud be :- Area of sector = \(\large \frac{1}{ 2} r^2\theta = \frac{1}{ 2} 6^2\times 2 \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

is there ny need to cnvrt radian into degree?

OpenStudy (anonymous):

no there is no need to convert radians to degree here

OpenStudy (anonymous):

as ganeshie pointed out - find the circumference and area of the whole circle first and then using proportion find out how much area will the 12 cm arc sector make

OpenStudy (anonymous):

three water sprinklers are placed at A, E and D on the rectangular field ABCD, Ebeing the mid-point of AD. Each can throw water in a circular region upto a radius of 4.2m. if the lenth and bredth of the field is 40m and 26m respectively, how much area of field is watered by them?

ganeshie8 (ganeshie8):

hey please close this question and post a new question.. . that way ur question shows up in new questions and others get to see :)

ganeshie8 (ganeshie8):

first "Close the Question"

ganeshie8 (ganeshie8):

do u see "Close Question" in green

ganeshie8 (ganeshie8):

right below ur question posted area,

ganeshie8 (ganeshie8):

scroll up... all the way up...